OpenStudy (anonymous):

Proof, if A is a unitary matrix, then det(A)=1

OpenStudy (anonymous):

I tried to prove it to \[A\epsilon \mathbb{C}^{2x2}\] but I didn't know how to do generalize the idea to an nxn case.

OpenStudy (anonymous):

Are you sure it isn't |det(A)| = 1?

OpenStudy (anonymous):

Well, anyway, I don't think it's true unless |det(A)| = 1. Here's the proof for what I stated though: 1 = det(I) = det(AA*) = det(A)det(A*) = det(A)det(A)* = |det(A)|^2 So, |det(A)| = 1

OpenStudy (anonymous):

If det(A)=1 then det(A)=1... it's trivial. Anyway thank you, I found the same proof using some facts about determinants I found in wikipedia.

OpenStudy (anonymous):

|det(A)| = 1 => det(A) = 1 or -1. It isn't trivial. The absolute value is required.

OpenStudy (anonymous):

Sorry I misread.

OpenStudy (anonymous):

I misread 2 times, one was when I read "I don't think it's true unless |det(A)| = 1" And then when I read the exercise in my homework. It was |det(A)|, not det(A) :p

OpenStudy (anonymous):

Ah. Good. Just checking. :)

OpenStudy (anonymous):

I'm closing this.